The ratio of spaniels to poodles at the dog show was 5 to 7.If there were 35 spaniels at the dog show, how many poodles were there?(the possibilities are.........35, 49, 25, or none of these)
Answer by askmemath(368)

Let the total number of dogs be = D
IF the ratio is 5:7 then the number of spaniels = 5/(5+7) x D = 35
So D = 84
Now the number of poodles = 7/(5+7) x 84 = 49